OpenStudy (anonymous):

(3sqrt7-9sqrt3)(4sqrt7+12sqrt3) multiply

OpenStudy (anonymous):

\[(3\sqrt{7}-9\sqrt{3})(4\sqrt{7}+12\sqrt{3})\]